# Understanding Dynamic Diffusion Process of LLM-based Agents under Information Asymmetry

### v2

@@ Abstract
Large language models have been used to simulate human society using multi-agent systems. Most current social simulation research emphasizes interactive behaviors in fixed environments, ignoring information opacity, relationship variability and diffusion diversity. In this paper, we study the dynamics of information diffusion in 12 asymmetric open environments defined by information content and distribution mechanisms. We first present a general framework to capture the features of information diffusion. Then, we designed a dynamic attention mechanism to help agents allocate attention to different information, addressing the limitations of LLM-based attention. Agents start by responding to external information stimuli within a five-agent group, increasing group size and forming information circles while developing relationships and sharing information. Additionally, we observe the emergence of information cocoons, the evolution of information gaps, and the accumulation of social capital, which are closely linked to psychological, sociological, and communication theories.
@@ 1 Introduction
To achieve this objective, we first introduce a two-tier general simulation framework designed to capture dynamic information diffusion. We also propose an agent attention mechanism Baars (2005); Chen (2016) that prioritizes critical information in a manner analogous to human information processing, enabling agents to handle multiple sources of information concurrently. We then examine the behaviors of agents under various external stimuli. Our study incorporates both macro-level and micro-level analyses to elucidate differences in information gaps, communication dynamics, and the evolving structure of information circles. We also explore the impact of integrating new agents and investigate various phenomena pertinent to psychology, communication, and sociology.
@@ 2.1 General Simulation Framework
Action  At each time step $T=\left\{t_{1},...,t_{10}\right\}$, we have agents $A=\left\{a_{1},...,a_{n}\right\}$. At the beginning of the simulation time step $t_{1}$, $n=5$. At each time step $t$ after this, n may increase based on the actions of each agent, up to a maximum of 5 per round. Each agent $a_{i}$ has profile $p_{i}$, relationship $r_{i}$, output action $o_{i}$, information diffusion $d_{i}$. At time step $t_{i}$, $p_{i}$ remains unchanged, $r_{i}$ has scale $r_{i}\in\left\{\text{positive},\text{negative},\text{general}\right\}$ with other agents. $o_{i}$ can be True or False and consists of two parts: changing the relationship $r_{i}$ and transmitting information $d_{i}$. The agent can independently choose to pass information to any agent in the current environment or to a new agent it defines itself. Therefore, the agent’s action decision-making must balance the initial information with other information, including discussions caused by profile similarities. After each action round, the environment updates the state of each agent based on the agent’s actions $O=\left\{o_{1},...,o_{n}\right\}$. This includes updating $r_{i}$ (subjective relationship) in the database, adjusting $d_{i}$ to reflect the corresponding receiver’s received_messages, and refreshing the agent’s actions for this round. After that, the environment updates the list of the latest agents and performs attention calculations as algorithm 1 and action decisions for the next round of agents.
2.2 Agent Construction
In this section, we introduce how to construct an agent. Especially, we propose a novel agent memory mechanism, referred to as the dynamic attention mechanism. This approach is motivated by the observation that, within our simulation, agents receive information from multiple sources during each interaction round, substantially increasing the total context length. Consequently, agents may struggle to discern which pieces of information warrant the most attention.
@@ 2.2 Agent Construction
In preliminary experiments, we employed a generic LLM-based agent that retained all information received across multiple rounds in its memory, making decisions based on this complete dataset. We observed that, under this design, a single agent’s actions across communication rounds remained highly similar, resulting in minimal active changes to interpersonal relationships throughout the simulation. Such uniformity in behavior diverges from patterns typically observed in human interactions Gong et al. (2023); Baqir et al. (2025); Stein and Harper (2012).
mean $\uparrow$ min max SD $\downarrow$
Generic Agent 0.80 0.63 0.94 0.08
Agent + Dynamic Attention 0.57 0.27 0.90 0.17
Table 1: Preliminary study for agent attention algorithm.
In this preliminary experiment, we conducted 24 simulation runs across eight distinct scenarios. As illustrated in the first row of Table 1, we present the outcomes for a generic LLM-based agent in terms of mean, minimum, maximum, and standard deviation (SD) of the similarity between actions in consecutive rounds. These similarity values were calculated by applying cosine similarity to action embeddings generated using Sentence-BERT Reimers and Gurevych (2019). Notably, the mean similarity reaches as high as 0.80, while even the least similar pair of actions exhibits a cosine similarity of 0.63. Furthermore, the small standard deviation indicates that this tendency is highly consistent across simulations, suggesting that the agents frequently repeated the same information over the course of the simulation.
@@
This observation stands in stark contrast to real-world information diffusion processes Li et al. (2017); Guille et al. (2013). Consequently, relying solely on an LLM’s intrinsic attention mechanisms over an extended context constrains the representation of how various pieces of information compete for an agent’s focus. Agents need more factors related to the real world (such as interpersonal relationships, information complexity, information changes) to assist them in making wise action decisions. To address these shortcomings, we propose an agent attention algorithm designed to mitigate these issues.
Agent with Dynamic Attention
@@ 7       $dict[s]\leftarrow(w,m)$;
[…]
$H(X)=-\sum_{i=1}^{n}(p_{i}+\epsilon)\log(p_{i}+\epsilon),\quad\epsilon=10^{-9}$ (1)
Figure 2: Information Gap (the blue bars) and Diffusion Gap (the red bars) for 12 asymmetric environments on four initial settings. Each simulation contains these two values. Differences between the two values represents the Diffusion Conversion Gap. The smaller the gap, the more individuals with known information tend to spread it, which means that the diffusion chain is relatively complete.
3 Experiment and Analysis
This section outlines the experimental design, data analysis, and findings. We examine how information asymmetry evolves within a multi-agent system by exploring various forms of asymmetric information and comparing the observed outcomes to established social science literature. This comparison demonstrates the viability of using LLM-based agents to simulate environments characterized by asymmetric information. Furthermore, we highlight novel insights made possible by multi-agent simulations—insights that are difficult to capture in traditional social science and communication studies, thereby illustrating new avenues for employing LLM-based agents in social science research.
@@ 3.1 Experimental Settings
To examine the formation of a dynamic information gap during agent information diffusion in an asymmetric open environment, we conducted a simulation and tested it in 12 different asymmetrical information environments. The simulation is developed based on the SOTOPIA Zhou et al. (2023) library and employs the GPT-4o mini model openai (2024) for the agent’s decision-making process. We randomly selected 5 agents from the 25 agents in Stanford Town Park et al. (2023) as the initial state group. Their profiles include gender, age, innateness, and occupation, and are evenly distributed. The group settings include the group’s topology and initial relationship.
@@
We jointly build an information asymmetry environment through information content and distribution mechanism. The main difference in the information content lies in its relevance to initial agents, and the distribution mechanism mainly affects the asymmetry generated directly at the source of information. Based on the Construal Level Theory Trope and Liberman (2010) in social psychology, we define five types of information content: other people’s gossip (OG), public policy (PP), legal cases (LC), and stakeholder (SH). Furthermore, we define three distribution mechanisms: information broadcast (BC), information unicast (OA), and broadcast by round (BCR), creating asymmetry at the source of information. BC represents the process of send the information to all five agents at the first round, while OA means only send information to one agent (agent 2 as the center in wheel and common node in circle). BCR means send information to one agent each round until the initial five agents know the information. Our information is generated by GPT-4o mini and is approximately 50 words long, as shown in table 6.
@@
We ran simulations three times for each topology corresponding to the information content and asymmetric mechanism, with the initial relationships between agents set to all positive or all negative.
3.2 Macro-level Analysis
We compare the information diffusion process of agent groups in different information asymmetric environments through the information gap, the diffusion gap, and information retention.
The information gap refers to the percentage of agents aware of the initial information compared to all agents. The diffusion gap indicates the percentage of agents who have shared the initial information within the group. Additionally, information retention measures how many rounds the initial information is maintained during the group’s diffusion process. Since the message sent by an agent is a short sentence rather than a long text, we use the Sentence-BERT method Reimers and Gurevych (2019) to compare the similarity between the message and the initial information. If the similarity is greater than 80%, it means the receiver is aware of the initial information, and the sender has successfully shared it. It is important to note that an individual can only share information they already know, simulating the natural process of people sharing thoughts or observations they have recognized. The results are summarized as follows.
Figure 3: Information retention for 12 asymmetric environments under four initial settings. The larger the value, the longer the initial information is retained in the diffusion process and the harder it is to be submerged.
@@ 3.2 Macro-level Analysis
Agenda-Setting Theory McCombs et al. (2013): Distributing information over time helps maintain relevant knowledge within a group, but is not effective for widespread sharing.  As shown in Figure 3, when information related to the initial five individuals (except the OG) is spread using BC and BCR, information retention is consistently higher with BC than with BCR. However, Figure 2 shows that the diffusion gap values for BCR are always smaller than those for BC. Although the initial five agents can know the initial information through BC and BCR, BCR delays the time. These two methods directly reflect the information source’s control over the "agenda." The sooner information is introduced to the group (as in BC), the more likely it is to capture individuals’ attention. It can also be concluded that in the information diffusion process of LLM agents, slowing down the time it takes for agents to access information can reduce large-scale dissemination, thereby increasing information retention within the group.
Social Identity Theory Brown (2000): Individuals tend to spread highly relevant content.  As shown in Figure 2, the largest difference between the information gap and the diffusion gap is concentrated in the OG information. This suggests that OG information is difficult to spread effectively within the current group. Individuals are more likely to share information that is highly relevant to themselves, whether for cooperation, understanding, in-depth discussion, or other purposes.
3.3 Micro-level Analysis
Figure 4: In this Social Capital Theory case study, agents 1, 3, 4, and 5 form relationships with new agents, creating distinct information circles within the growing group. Nodes represent individual agents, with colors indicating their lineage (the agent and its new recruits are in the same lineage). Node distance and edge color depth reflect the similarity between the current message and those of all agents. Agents who did not take action are not recorded.
In this section, we focus on the agent’s action to understand the information diffusion in an open environment with information asymmetry.
Social Behavior in information diffusion  First, we observed that agents have social motivations when spreading information (shown in Table 2), such as seeking cooperation, altruism, support, or discussion. When the initial information is sufficient to prompt an agent to spread it, the agent may seek cooperation based on this information. However, if the initial information is insufficient to influence the agent’s spreading behavior, the agent may choose not to act in that round or may communicate based on other received information and the existing profiles of agents. This behavior reflects that, regardless of whether the content of the communication is related to the initial information, similar social motivations will still arise.
(a) positive 46% negative 54%
(b) circle 50% wheel 50%
(c) BC 28% OA 42% BCR 30%
(d) OG 24% PP 26% LC 24% SH 26%
(e) OG 0% PP 12.5% LC 0% SH 87.5%
Table 3: The first four rows illustrate how different information asymmetry factors affect new agent diffusion. (a) and (b) represent the initial agent group settings, (c) and (d) show the external information asymmetry environment, and (e) is the proportion that new agents receive over 80% similarity with the initial information.
agent id total received average similarity $\uparrow$ min max standard deviation $\downarrow$
2 23 0.41 0 0.99 0.24
3 10 0.67 0.02 0.99 0.33
6 10 0.82 0.54 1 0.12
1 5 0.61 0.30 0.89 0.23
5 4 0.59 0.42 0.92 0.16
7 3 0.74 0.67 0.81 0.06
4 2 0.71 0.71 0.70 0
8 2 0.82 0.82 0.82 0
Table 4: Case study on Information Cocoon: In a simulation, higher pairwise similarity and lower standard deviation of agents’ received information correlate with stronger cocoon formation.
Diffusion in Open Environment  The open environment enables agents to recruit new members in any round, with the profile of each new member determined by the recruiting agent. We summarized the results of 48 simulations, involving 304 instances of agents recruiting new members, as shown in Table 3. The analysis examined the effects of two initial settings (relationship, topology) and two factors related to information asymmetry (information content, distribution mechanism) on the number of new agents recruited. When the relationship is negative or when the outside world releases information in the form of OA, agents are more inclined to send message to new agents. Other conditions have little effect on the number of new agents. However, as seen in the proportions in (d) and (e), even though the content of the information has little effect on an agent’s behavior in recruiting new members, it significantly influences whether the agent shares the initial information with the new recruit. When the information is public and involves clear interests, agents are more likely to communicate it to new recruits. However, if the information pertains to private matters, it is not shared externally but discussed within the group.
Social Capital Theory Lin et al. (2001)  According to Figure 4, we found that agent 4 continuously expanded its information circle in the group by establishing connections with new agents, and at the same time won the attention of agent 2. This illustrates how agents gather more information resources through social networks. Gaining more public attention makes it easier to become an opinion leader in the group.
Information Cocoon Yuan and Wang (2022)  Table 4 shows that in this simulation, agent 6 received a total of 10 messages, with a pairwise similarity of 0.82 and a low standard deviation. This shows that even though agent 6 received a lot of information, it formed an information cocoon, making it more difficult to make decisions about dissemination actions based on diverse information.

### v3

# Attention Mechanism for LLM-based Agents Dynamic Diffusion under Information Asymmetry

@@ Abstract
Large language models have been used to simulate human society using multi-agent systems. Most current social simulation research emphasizes interactive behaviors in fixed environments, ignoring information opacity, relationship variability, and diffusion diversity. In this paper, we first propose a general framework for exploring multi-agent information diffusion. We identified LLMs’ deficiency in the perception and utilization of social relationships, as well as diverse actions. Then, we designed a dynamic attention mechanism to help agents allocate attention to different information, addressing the limitations of the LLM attention mechanism. Agents start by responding to external information stimuli within a five-agent group, increasing group size and forming information circles while developing relationships and sharing information. Additionally, we explore the information diffusion features in the asymmetric open environment by observing the evolution of information gaps, diffusion patterns, and the accumulation of social capital, which are closely linked to psychological, sociological, and communication theories.
@@ 1 Introduction
To achieve this objective, we first introduce a two-tier general simulation framework designed to capture dynamic information diffusion. We also propose an agent attention mechanism Baars (2005); Chen (2016) that prioritizes critical information in a manner analogous to human information processing, enabling agents to handle multiple sources of information concurrently. We then examine the behaviors of agents under various external stimuli. The Dynamic Attention algorithm helps to reduce the Action Similarity Bias and increase the Relationship Perception Frequency, thereby proving the effectiveness of the algorithm in assisting agents in handling complex social scenarios in social simulations. Meanwhile, we explored the social phenomenon of multi-agent information diffusion in the asymmetric open environment.
@@ 2.1 General Simulation Framework
Generic LLM Agent Dynamic Attention Agent
SH-BC-negative OG-BCR-circle SH-BC-negative OG-BCR-circle
circle wheel positive negative circle wheel positive negative
Action Similarity Bias $\downarrow$
GPT-4o-mini 0.99 0.99 0.95 0.92 0.90 0.90 0.86 0.87
llama-3.3-70b 0.96 0.96 0.94 0.94 0.89 0.89 0.87 0.88
Relationship Perception Frequency $\uparrow$
GPT-4o-mini 0.67 0 0.67 2 2 2.67 2 5.33
llama-3.3-70b 0 0 0 0 0.33 0.33 0 0.33
Table 1: Preliminary test for LLM’s deficiency on information diffusion in social simulations. The table shows the mean value for three times simulation for every setting.
@@
Action  At each time step $T=\left\{t_{1},...,t_{10}\right\}$, we have agents $A=\left\{a_{1},...,a_{n}\right\}$. At the beginning of the simulation time step $t_{1}$, $n=5$. At each time step $t$ after this, n may increase based on the actions of each agent, up to a maximum of 5 per round. Each agent $a_{i}$ has profile $p_{i}$, relationship $r_{i}$, output action $o_{i}$, information diffusion $d_{i}$. At time step $t_{i}$, $p_{i}$ remains unchanged, $r_{i}$ has scale $r_{i}\in\left\{\text{positive},\text{negative},\text{general}\right\}$ with other agents. $o_{i}$ can be True or False and consists of two parts: changing the relationship $r_{i}$ and transmitting information $d_{i}$. The agent can independently choose to pass information to any agent in the current environment or to a new agent it defines itself. Therefore, the agent’s action decision-making must balance the initial information with other information, including discussions caused by profile similarities. After each action round, the environment updates the state of each agent based on the agent’s actions $O=\left\{o_{1},...,o_{n}\right\}$. This includes updating $r_{i}$ (subjective relationship) in the database, adjusting $d_{i}$ to reflect the corresponding receiver’s received_messages, and refreshing the agent’s actions for this round. After that, the environment updates the list of the latest agents and performs attention calculations as algorithm 1 and action decisions for the next round of agents. More detailed explanations are available in Appendix C.
2.2 LLM’s Deficiency in Complex Social Environments
After building the general simulation framework, we set up preliminary tests for learning generic LLM competency. We employed a generic LLM-based agent that retained all information received across multiple rounds in its memory, making decisions based on this complete dataset. We find two kinds of deficiency of generic LLM on information diffusion in social simulations. The test experiment settings include four kinds of information asymmetry environments, and the detailed information list in Appendix D (all the simulations ran by SOTA models GPT-4o-mini-2024-07-18 and llama-3.3-70b-instruct).
LLM outputs similar actions in different rounds.
@@ 2.2 LLM’s Deficiency in Complex Social Environments
We observed that, under this design, a single agent’s actions across communication rounds remained highly similar throughout the simulation. This indicates that agents relying on generic LLMs have difficulty generating diverse and meaningful information dissemination behaviors. Such uniformity in behavior diverges from patterns typically observed in human interactions Gong et al. (2023); Baqir et al. (2025); Stein and Harper (2012). In contrast, Dynamic Attention helps the agent reduce the likelihood of outputting similar messages in multiple rounds, thereby increasing the diversity of the messages.
LLM lacks perception of social relationships.
@@
Another finding focuses on how the LLM agent uses the relationship and how the relationship changes during the simulation. We observe LLM remains a low relationship perception frequency in different asymmetry environments, which means the social relationship cannot effectively influence LLM’s decision and is also not included in the decisions. This observation stands in stark contrast to real-world information diffusion processes Li et al. (2017); Guille et al. (2013). Consequently, relying solely on an LLM’s intrinsic attention mechanisms over an extended context constrains the representation of how various pieces of information (sent by other agents with different social relationships) compete for an agent’s focus. Agents need more factors related to the real world (such as interpersonal relationships, information complexity, and information changes) to assist them in making wise action decisions. To address these shortcomings, we propose an agent attention algorithm designed to mitigate these issues.
2.3 Agent with Dynamic Attention
@@ 7       $dict[s]\leftarrow(w,m)$;
Subsequently, the agent refines these weights by comparing short-term memory with long-term memory. This step emulates the process by which humans recall information sources and consider past exchanges. To highlight messages that exhibit the greatest level of transformation during transmission, the algorithm calculates the change in the entropy value of the corresponding information source from the previous round to the current round. For each round, if the message output by the agent is $p$, then calculate the entropy value of this agent in round $i$ according to Equation 1. Lastly, the algorithm further increases the weight of messages originating from agents with whom there have been the most frequent interactions, as inferred from past actions.
$H(X)=-p_{i}\log(p_{i})$ (1)
3 Experiment and Evaluation
This section outlines our experiment design and the ablation study for Dynamic Attention Algorithm. We establish 12 types of asymmetric environments based on information diffusion and test the performance of the SOTA LLM using the Dynamic Attention algorithm. In the ablation experiment, we compare with generic llm (with all memory) and generic llm (with last round memory), and verified the effectiveness of the algorithm from two perspectives: the agents’ behavior of and the utilization of social relationships.
@@ 3.1 Experimental Settings
To examine the formation of agent information diffusion in an asymmetric open environment, we conduct a simulation and tested it in 12 different asymmetrical information environments. The simulation is developed based on the SOTOPIA Zhou et al. (2023) library, Redis database Redis , and employs the SOTA model GPT-4o-mini and llama-3.3-70b-instruct openai (2024); huggingface for the agent’s decision-making process. We randomly select 5 agents from the 25 agents in Stanford Town Park et al. (2023) as the initial state group. Their profiles include gender, age, innateness, and occupation, and are evenly distributed. The group settings include the group’s topology and initial relationship.
@@
We jointly build an information asymmetry environment through information content and distribution mechanism. The main difference in the information content lies in its relevance to initial agents, and the distribution mechanism mainly affects the asymmetry generated directly at the source of information. Based on the Construal Level Theory Trope and Liberman (2010) in social psychology, we define four types of information content: other people’s gossip (OG), public policy (PP), legal cases (LC), and stakeholder information (SH). Furthermore, we define three distribution mechanisms: information broadcast (BC), information unicast (OA), and broadcast by round (BCR), creating asymmetry at the source of information. BC represents the process of send the information to all five agents at the first round, while OA means only send information to one agent (agent 2 as the center in wheel and common node in circle). BCR means send information to one agent each round until the initial five agents know the information. Our information is generated by GPT-4o-mini and is about 50 words long, as shown in table 6.
@@
We run simulations three times for each topology corresponding to the information content and asymmetric mechanism, with the initial relationships between agents set to all positive or all negative.
3.2 Ablation Study
In this section, we compare three scenarios: the dynamic attention algorithm (Dynamic Attention), the generic LLM with all-round memory (Generic LLM-all), and the generic LLM with last-round memory (Generic LLM-last). Based on the experiment settings, we choose different asymmetry environments to verify the dynamic attention algorithm. We list the results in Appendix E.
4 Result and Analysis
RQ 1: How do asymmetry settings influence agents’ information diffusion?
Distributing information over time helps maintain relevant knowledge within a group, but it is not effective for widespread sharing.
We calculate the Information Retention for every simulation, and the result shows simulations using the BCR mechanism remain a 3.14 rounds for the initial information retention in the group diffusion process, while in the BC setting, the value is only 2.77. As shown in Figure 2, in different models, we can observe that although the BCR mechanism keeps the initial information in the population for a longer period compared to the BC mechanism, it does not lead to more spread within the population. This phenomenon is in line with the Agenda-Setting Theory in communication studies. When the source of information uses different methods to disseminate the initial information, the group will exhibit corresponding different behaviors.
Different LLMs may vary in their willingness to spread information, but their performance trends are consistent in different environments of information asymmetry.
We observe similar trends in the differences among the three information distribution mechanisms. This indicates that the information gap and diffusion gap reach their highest levels when the information source broadcasts. Following this is BCR, and then comes OA. We can also observe that the llama-3.3-70b-instruct model consistently shows differences in its willingness to disseminate information content across various initial group relationships.
The agent tends to disseminate internally the information that they are the party involved, while spreading externally information that is beneficial to the public.
Although the gap in the diffusion intentions presented by the gpt-4o-mini model is relatively small, it can still be observed that both models are more inclined to disseminate information that is related to their interests or where they are the parties involved. More importantly, based on the values of LC and SH in Table 2, when the agent is the party involved in the event, they tend to have internal discussions rather than disseminate the information to other agents outside. The difference is that the SH information not only includes the initial five aspects of the agent’s interests, but also contains beneficial information for the public. In such cases, the agent is more likely to disseminate this information to the outside world.
Figure 2: Information Gap (the grey bars) and Diffusion Gap (the green bars) for 12 asymmetric environments on four initial settings. Each simulation contains these two values. Differences between the two values represent the Diffusion Conversion Gap. The smaller the gap, the more individuals with known information tend to spread it, which means that the diffusion chain is relatively complete.
RQ 2: How does social relationship and topology change after the agent’s information diffusion in "open environment"?
In our design, the agent can choose to send messages to any new agent (which does not exist in the environment) at any round. But this does not mean that they will choose to pass on the relevant content of the initial information to the new agent. Therefore, we counted the number of new agents in different information environments, as well as the similarity between the information received by the new agents upon their initial entry into the group and the initial information (shown in Table 2). In the stakeholder information environment, agents are most inclined to communicate with new agents, which is the opposite of the legal case scenario. Similarly, in the stakeholder environment, the content exchanged between the agent and the new agent is the most similar to the initial information.. Furthermore, all the cases where the new agent received information that was 80% similar to the initial information were concentrated in the Stakeholder information.
OG PP LC SH
the average number for agent increase 1.196 1.167 0.944 1.375
new agent & receive initial information 0.293 0.267 0.304 0.404
Table 2: Analysis for new agents in the asymmetric open environment
In the same experimental simulation setting, different models perform variably, which contrasts with common perceptions. We calculate the changes made by the model to the social relationship and topology (represented by the agent number). In Table 3, the ratio of actions that changed the existing relationships or added new agents to the total number of simulations is shown for all simulations under the corresponding environmental settings. gpt-4o-mini tends to directly pull new agents into the group, while llama-3.3-70b-instruct usually makes changes within the existing relationships.
gpt-4o-mini llama-3.3- 70b-instruct
BC OA BCR BC OA BCR
change relationship / / 0.08 1 1.08 1.17
add new agent 1.42 1.25 1.33 0.08 / 0.17
Table 3: Comparison for different LLM social behavior in "asymmetric open environment".
Figure 3: In this Social Capital Theory case study, agents 1, 3, 4, and 5 form relationships with new agents, creating distinct information circles within the growing group. Nodes represent individual agents, with colors indicating their lineage (the agent and its recruits are in the same lineage). Node distance and edge color depth reflect the similarity between the current message and those of all agents. Agents who did not take action are not recorded.
@@ 4 Result and Analysis
RQ 3: What’s the relation between received messages and the agent’s social behavior?
In the experiments, we observe the social characteristics displayed by the agent when conveying information. For example, after discussing activity details, the agent may express a desire to cooperate or exhibit altruistic behavior by bringing other relevant agents into the group and relaying pertinent information. Moreover, regardless of whether the information sent by the agent is related to the initial information, it will always reflect social features. Specific instances are presented in Table 4.
On the other side, different weights are assigned to the received messages to influence the agent’s behavior by competing for its attention. The experimental results show that obtaining a high-weight message does not necessarily directly prompt the agent to spread this message, nor does it directly lead the agent to converse with the sender of the high-weight message in the next round. Due to the agent’s influence or other factors, messages related to the initial information may not receive a high weight. This also explains why initial environmental information might cease to be disseminated in a certain round. This setup integrates the changes in social relationships and the messages themselves over multiple rounds into the agent’s information diffusion process, guiding the LLM to take actions that better suit the agent’s current situation. Detailed cases are provided in Appendix B.
Overall, we observe many cases of Social Capital Theory in different simulations. Although in our research, agents don’t have economic settings, they also accumulate information capital by sending messages to other agents, pulling different new agents into the group, and building wise social relationships with initial and subsequent agents.